The interest rate for the first five years of an $100,000 mortgage loan is 9.4% compounded semiannually. Monthly payments are calculated using a 20-year amortization.

a.

What will be the principal balance at the end of the five-year term? (Do not round intermediate calculations and round your final answer to 2 decimal places.)

  Principal balance $

    

b.

What will be the monthly payments if the loan is renewed at 6.8% compounded semiannually (and the original amortization period is continued)? (Do not round intermediate calculations and round your final answer to 2 decimal places.)

  Payment $ per month
Answer #1

1.
=FV((1+9.4%/2)^(2/12)-1,12*5,PMT((1+9.4%/2)^(2/12)-1,12*20,-100000),-100000)=88956.5129688202

2.
=PMT((1+6.8%/2)^(2/12)-1,12*15,-FV((1+9.4%/2)^(2/12)-1,12*5,PMT((1+9.4%/2)^(2/12)-1,12*20,-100000),-100000))=784.997494737316